Minute of KAGRA ERun1 summary meeting
- Date: 2019/6/11 15:00~16:00
- Attendance: Miyakawa, Aso, Oshino, Ogaki, Kokeyama, Nakano, Enomoto, Kozakai, Miyoki, Miyo, Yokozawa, Haino, Suzuki, Koyama, Nakagawa, Tanaka, Sugimoto, Stefan
- Zoom: Yokoyama, Youngmin, Inoue, Takahashi, Shoda, Uchiyama, Takeda, Akutsu, Yamada, Suzuki, Sawada
Over all summary
- X arm lock
- Lock was quite stable.
- PR3 troubles for the first 30 min, probably due to temperature.
- Oplev itself is something wrong, and PR3 is largest one.
- PMC drift due to the temperature change, and out of the range of PZT lost lock.
- Green power was lower now, but not on Saturday, but on this week.
- First Sensitivity monitor
Parallel works during Run
- ETMY installation work
- DAQ trouble
- Lock flag, Science mode, effective data
- Duty cycle
- Data rate
- Yokozawa will collect necessary data.
- Long term temperature investigation
- Calibration for Mass lock
- Limited by ADC noise. Needs whitening filters.
- Noise hunting for frequency noise
- Needs MICH calibration.
- Gardian for DAQ
- Real time camera
- COM team should take a rest during Run.
- Needs shift ASAP
- With LIGO commissioners.
- Saturday or Friday?
- 7/12 or 13.
- ITMY trouble?
Minute of temperature control meeting
- Date: 2019/6/11 12:15~13:00
- Attendance: Miyakawa, Yokozawa, Enomoto, Nakano, Kokeyama, Miyoki, Aso, Sugimoto, Stefan
- Cause of instability in PSL.
- Sensor position.
- Not circulating well.
- To do
- Stopping Water cooler around MCE.
- Move sensor position.
Stopping all AC in the PSL? -> temperature goes to more than 30deg.